How many milliliters of a 0.83 M calcium hydroxide solution are required to completely react with 407 mL of a 0.31 M nitric acid solution? Hint: This means remove all of the protons form the acid.

Answer with no decimal places.
Answer #1

Balanced chemical equation is:

2 HNO3 + Ca(OH)2 ---> Ca(NO3)2 + 2 H2O

Here:

M(HNO3)=0.31 M

M(Ca(OH)2)=0.83 M

V(HNO3)=407.0 mL

According to balanced reaction:

1*number of mol of HNO3 =2*number of mol of Ca(OH)2

1*M(HNO3)*V(HNO3) =2*M(Ca(OH)2)*V(Ca(OH)2)

1*0.31 M *407.0 mL = 2*0.83M *V(Ca(OH)2)

V(Ca(OH)2) = 76.006 mL

Answer: 76.0 mL